OpenStudy (anonymous):

use the formula d=rt find the value of the missing variable. d=160 miles,r=32 mph

OpenStudy (whpalmer4):

\[d = rt\] We know \(d = 160\text{ miles}\) and \(r = 32\text{ miles/hour}\) so \(t\) is the missing variable. To solve for \(t\), we divide both sides of the equation by \(r\), giving us: \[\frac{d}{r} = \frac{rt}{r} \]\[t = \frac{d}{r}\]Now plug in the known values of \(d\) and \(r\) to find the value of \(t\).

OpenStudy (anonymous):

2

OpenStudy (whpalmer4):

how on earth did you come up with 2?

OpenStudy (anonymous):

ok im damb

OpenStudy (whpalmer4):

do you drive?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

\[t = \frac{d}{r} = \frac{160}{32}=\]Can you divide 160 by 32?

OpenStudy (whpalmer4):

What does 32+32= What does 32+32+32= What does 32+32+32+32= What does 32+32+32+32+32= That should give you a hint about the answer...

OpenStudy (anonymous):

160

OpenStudy (whpalmer4):

no. 160/32 = ? how many 32s did you have to add up to get 160?

OpenStudy (anonymous):

5

OpenStudy (whpalmer4):

okay. that means that 32*5 = 160 and 160/32 = 5.

OpenStudy (anonymous):

ok